# From Single Genes to Populations: Quantifying Broken Detailed Balance in Transcription

### v2

@@ III Entropy production of transcription in a population of cells
The results of the extrinsic noise analysis are shown in Fig. 6. The line plots in Figs. 6(a), (c) and (e) show how the relative error changes as a function of the variance of the extrinsic noise for fixed mean, while the density plots show the errors for given parameter sets at fixed $\overline{\theta}=\mathrm{Var}(\theta)$. Generally, it is found that having cell-to-cell variability in the population increases $\dot{s}_{\mathrm{mes}}$, and that in general the system is further away from detailed balance—see Fig. 6(a), (b), (e) and (f). The exception to this rule is that when the noise is on $\sigma_{\mathrm{off}}$, $e_{\sigma_{\mathrm{off}}}$ can be positive or negative depending on the values of $\sigma_{\mathrm{on}}$ and $d$—see Fig. 6(c) and (d). The result is that the EPR of a system with extrinsically varying kinetic parameters in a population does not correspond with the “representative cell” comprising the mean of all kinetic parameters.
@@
Perhaps the most interesting results come from $e_{\sigma_{\mathrm{off}}}$ ($e_{\sigma_{\mathrm{on}}}$), where not only can $\sigma_{\mathrm{off}}$ ($\sigma_{\mathrm{on}}$) variability increase $\dot{s}_{\mathrm{mes}}$ for $\overline{\sigma}_{\mathrm{off}}\ll\sigma_{\mathrm{on}}$ ($\sigma_{\mathrm{off}}\gg\overline{\sigma}_{\mathrm{on}}$), but also significantly lower it for most of the space of parameters outside that regime (see Fig. 6(d)). Additionally, in Fig. 6(c) it is shown that $e_{\sigma_{\mathrm{off}}}$ ($e_{\sigma_{\mathrm{on}}}$) is non-monotonic in the variance of the extrinsic noise, and that for small degradation rates there is a finite valued maximum. The summary is that the irreversibility inferred in a single cell can be significantly perturbed by the presence of cell-to-cell variability. This leads to the surprising conclusion that the suppression of cell-to-cell variability is may be more beneficial on some parameters than for others, that is, if the results of the previous section on the observed avoidance of high $\dot{s}_{\mathrm{mes}}$ (i.e., mesoscopic energy expenditure minimization) are to be taken seriously.
IV Conclusions
@@
This study highlights the role of detailed balance and irreversibility in the canonical two-state model of mRNA expression. This includes: the nature of probability fluxes in the steady state (Fig. 3); the analytics of entropy production for a single gene; detailed data analyses of seven kinetic parameter datasets (Figs. 4 and 5 and corresponding figures in the SI); and the counterintuitive effects of cell-to-cell variability on the entropy production rate across a population of cells (Fig. 6). Our results have highlighted a potential mesoscopic analog of the principle of energy expenditure minimization that reveals itself through the EPR of mRNA expression [72]. Notably, this is not a thermodynamic phenomenon—we have shown that the lower bound on the EPR predicted by our mesoscopic analysis is not tight compared to order of magnitude estimates. Finally, our results highlight the benefits of transcriptional bursting from the perspective that a burstier behaviors can lead to a lower EPR (Fig. 5(d)).
@@
One of the limitations of our study is that we consider only a two-state gene model of transcription, whereas intricate experiments have shown that transcription in higher-order organisms is best modeled by more than two gene states [21, 40, 41, 16, 42, 43, 44]. Recent experimental data even reveals that some genes have sub-Poissonian mRNA expression in fission yeast [77], a behavior that aligns with earlier theoretical predictions [78, 79, 80]. Although the analytic framework introduced in this study can likely be extended to consider multiple gene states, there are no studies that infer transcriptional kinetics across thousands of genes for models more complex than a two-state model [45, 46, 19]. Until such datasets are widely accessible, analytic results for the EPR of more complex gene state models cannot be fully utilized. Finally, our framework assumes gamma-distributed extrinsic noise, but heavy-tailed variation (e.g., log-normal) may further alter EPR bounds. Future work could explore this numerically or via approximations, building on studies like [53].
[…]
@@
The finding that cell-to-cell variability can amplify or suppress irreversibility interacts with the seminal work of Lestas et al. [83] several ways. Therein, they investigated the limits of noise suppression given thermodynamic and information theoretic constraints, and derived quantitative bounds on the degree to which noise suppression is actually possible. Their findings, that the suppression of noise in one component of a system may amplify it in another, has clear parallels to the findings presented here, in that accepting a degree of cell-to-cell variability can lead to an ensemble of cells that is more reversible than an equivalent population of identical cells. Where noise is tolerated in the cellular environment is the result of a complex evolutionary process, but this process may not depend only on the properties of a single cell, but on the properties of a population of cells. Our study hints to further evolutionary constraints, that not only is the placement of noise and its suppression important in the context of intrinsic stochasticity, but in where extrinsic noise is tolerated, and in how it is suppressed.
@@
In future studies, it would be interesting to study the information theoretic properties of mRNA expression with more complex gene-state architectures—especially if advances in live-seq [84] and merFISH [85] data allow for comprehensive datasets of kinetic parameters for more microscopic models of mammalian transcription. Time-resolved mRNA sequencing would also allow for studies into the breaking of detailed balance in dynamic cellular environments, with fluctuating and correlated mRNA expression, using methods introduced in [5, 38]. Finally, an experimental understanding of how the kinetic parameters of the telegraph model fluctuate across a population of cells could give credence to some of the implications of cell-to-cell variability on reversibility found in this study, and of other studies related to the explicit modeling of extrinsic noise in addition to intrinsic stochasticity [24, 86, 52, 87]. How cells decide where to tolerate cell-to-cell variability is an important open problem.
@@
Acknowledgements: This work was supported by a Lou Schuyler grant from the Santa Fe Institute and National Science Foundation grants DMR-191073 and 2133863. I would like to thank Jacob Calvert, Christopher Lynn, Anish Pandya, Brandon Schlomann, Harrison Hartle, Artemy Kolchinsky and Ramon Grima for discussions and constructive feedback. Special thanks go to Ramon Grima and Artemy Kolchinsky for invaluable suggestions.
@@
Data availability: The data for telegraph model parameters used in this paper can be found in [19, 46, 45]. The data for the mRNA degradation rates can be found in [71].
@@
Author contributions: JH conceived, designed, implemented and wrote the study.
@@
In this section, we solve for the eigenvalues of the master equation describing the two-state gene model in Eq. (1). The reaction scheme with $\rho_{\mathrm{off}}=0$ has already been solved in time in the present form [48, 88] and additionally for the case where mRNA production is bursty [89]. Here we specify the calculation for non-zero $\rho_{\mathrm{off}}$. Physically, the eigenvalues of the master equation correspond to the inverse of the fundamental timescales governing the relaxation towards the steady state. We find the eigenvalues by imposing physical constraints on the generating function, essentially that the generating function’s power series is real and non-singular for $z\in[-1,1]$.
@@
Using the separation of variables ansatz $G(x,t)\sim e^{-\lambda_{m}t}f_{m}(x)$, which arises naturally from the linear structure of the master equation, leads to [59, 61],
@@
The second solution here is only linearly independent if the exponent is not a integer less than or equal to 0. Now, the condition on $f_{m}(z)$ here necessary to determine the $\lambda_{m}$ is that the powers of $(z-1)$ pre-multiplying the confluent hypergeometric functions must be integer powers. This is standard practice in eigenfunction solutions which are hypergeometrics, and it means that each $f_{m}(z)$ is real and non-singular for $z\in[-1,1]$, and corresponds to the infinite state limit of other similar solutions found in the literature for Moran-like processes [90, 91, 92]. Additionally, the exponential pre-factor and the hypergeometric functions themselves are already real and non-singular for finite $z$. Enforcing that $\lambda_{m}=dm$ and $\lambda_{m}-\Sigma=dm$ for integer $m\in\{1,2,3,\ldots\}$ then gives two respective sets of eigenvalues, $\lambda_{m}^{1}=dm$ and $\lambda_{m}^{2}=dm+\Sigma$ that are dependent on the rates of gene switching compared to the degradation rates of the mRNAs.
@@
These eigenvalues represent the inverse of the relaxation time scales of the system, and are in correspondence with results in the fast-switching (marginal single gene state) limit found in [93]. This almost completes the solution for the generating function of the telegraph model, aside from determination of the constants $C_{m}^{1}$ and $C_{m}^{2}$ which in principle can be done using methods related to Sturm-Liouville theory alongside knowledge of the initial conditions of $G$. For a relevant summary on Sturm-Liouville methods see [91, Appendix A], for the initial conditions of the telegraph model see the Appendix of [48].
S3 Calculating the entropy production rate
Following Schnakenberg [60, Eqs. (7.6)-(7.11)], the entropy production rate (EPR) is given by,
@@
The first term on the right-hand side is simply $\mathrm{d}S(t)/\mathrm{d}t$, where $S(t)=-\sum_{\mathbf{x}}P(\mathbf{x},t)\ln(P(\mathbf{x},t))$, i.e., the derivative of the entropy change of the system. The second term in the sum is then the entropy change due to coupling with external reservoirs, typically known as the entropy flow—the literature commonly defines this as the negative of the entropy flow rate, since one considers work done by the system, not on the system. In the case where the system is in a non-equilibrium steady state, then the entropy change in the system becomes zero, i.e., $\mathrm{d}S(t)/\mathrm{d}t=0$, and the EPR becomes the negative of the entropy flow rate [67, 27, 66],
@@
Once the connected states have been identified, we arrive at Eq. (9) in the main text. In Eq. (9) there are 2 sums that need to be evaluated. The first is $\sum_{n=1}^{\infty}a_{n}$ which simply evaluates to $-b_{1}$ since $a_{n}=b_{n+1}-b_{n}$. One also needs to evaluate the sum $\mathcal{S}=\sum_{n=1}^{\infty}b_{n}$, where, for reference, $b_{n}$ is given by
@@
S4 Extrinsic noise expressions
@@
S5 Marginality and coarse-graining of gene expression
A classic form of model reduction is that of timescale separation, a type of coarse-graining, which is often employed in gene expression under the ‘fast gene switching assumption’, i.e., the condition that $\mathrm{min}\{\sigma_{\mathrm{off}},\sigma_{\mathrm{on}}\}\gg\mathrm{max}\{\rho_{\mathrm{off}},\rho_{\mathrm{on}},d\}$ [94]. Although there are several ways of employing this via quasi equilibrium or steady state assumptions [95, 96, 97], the most principled way is via the method of ‘averaging’ [33, 98]. Experimental evidence for fast gene switching can be found in [99] among other studies. Denoting $f=\sigma_{\mathrm{off}}/(\sigma_{\mathrm{off}}+\sigma_{\mathrm{on}})$, the method of averaging then gives the following timescale reduced model valid under the fast gene switching assumption,
[…]
where $\rho=f\rho_{\mathrm{off}}+(1-f)\rho_{\mathrm{on}}$, which is nothing other than a one-dimensional microscopically reversible Markov chain which satisfies detailed balance and hence has zero entropy production in the steady state [100]. However, taking the same fast switching limit of Eq. (10) reveals,
which is generally non-zero. This trivial example clearly shows that although common methods of model reduction may capture the dynamics of stochastic gene expression, they lose thermodynamic information, a point made in numerous other studies [34, 69]. Notably, as stated in [69],
A more salient aspect of the analysis presented herein regards the interpretation of the marginal detailed balance without timescale separation arguments. Experiments often only give insight into the state of mRNA or protein numbers, and rarely into the gene states that are theoretically constructed. Hence, if all that can be observed is the mRNA dynamics, the two-state model is equivalent to a birth-death process with a time-dependent mRNA production rate $\rho(t)$, as shown in Fig. S1. Although the equivalent system has a noisy time-dependent rate (whose statistics are specified in the caption), it is still Markovian, and it provides an example of non-zero entropy production in a system that satisfies detailed balance. That is, a case of detailed balance with non-zero EPR. The caveat here is that although the dynamics are Markovian, the time-dependence of $\rho(t)$ allows for this seemingly contradictory situation. When $\rho_{\mathrm{off}}=0$ one can even have detailed balance and an infinite entropy production rate. This provides a striking departure from the standard narrative that detailed balance is reserved for systems in contact with a single reservoir of energy and/or particles or multi-component closed systems [58, 59]. Results in preparation seem to show that systems with generically connected gene states additionally share this “detailed-balance without equilibrium” property. A key takeaway is that although detailed-balance and zero EPR are often stated as being synonymous, this is not necessarily true when the time-independence of the kinetic rates is violated, as can be the case for systems with hidden Markov states.
@@
Figure S2: Analog of Fig. 4 but for 1,436 G2M cell-cycle stage mouse fibroblast genes [19].
Figure S3: Analog of Fig. 4 but for 5,623 mouse fibroblast genes from [46].
Figure S4: Analog of Fig. 4 but for 5,227 C57 mouse embryonic stem cell genes from [45].
Figure S5: Analog of Fig. 4 but for 4,925 CAST mouse embryonic stem cell genes from [45].
Figure S6: Analog of Fig. 4 but for 5,858 C57 mouse fibroblast genes from [45].
Figure S7: Analog of Fig. 4 but for 5,854 CAST mouse fibroblast genes from [45].
@@
Figure S9: Analog of Fig. 5 but for G2M cell-cycle stage mouse fibroblasts [19].
Figure S10: Analog of Fig. 5 but for mouse fibroblast cells from [46].
Figure S11: Analog of Fig. 5 but for C57 mouse embryonic stem cells from [45].
Figure S12: Analog of Fig. 5 but for CAST mouse embryonic stem cells from [45].
Figure S13: Analog of Fig. 5 but for C57 mouse fibroblast cells from [45].
Figure S14: Analog of Fig. 5 but for CAST mouse fibroblast cells from [45].

### v3

@@ Introduction
The results of the extrinsic noise analysis are shown in Fig. 8. The line plots in Figs. 8(a), (c) and (e) show how the relative error changes as a function of the variance of the extrinsic noise for fixed mean, while the density plots show the errors for given parameter sets at fixed $\overline{\theta}=\mathrm{Var}(\theta)$. Generally, it is found that having cell-to-cell variability in the population increases $\dot{s}_{\mathrm{mes}}$, and that in general the system is further away from detailed balance—see Fig. 8(a), (b), (e) and (f). The exception to this rule is that when the noise is on $\sigma_{\mathrm{off}}$, $e_{\sigma_{\mathrm{off}}}$ can be positive or negative depending on the values of $\sigma_{\mathrm{on}}$ and dd—see Fig. 8(c) and (d). The result is that the EPR of a system with extrinsically varying kinetic parameters in a population does not correspond with the “representative cell” comprising the mean of all kinetic parameters.
@@
Perhaps the most interesting results come from $e_{\sigma_{\mathrm{off}}}$ ($e_{\sigma_{\mathrm{on}}}$), where not only can $\sigma_{\mathrm{off}}$ ($\sigma_{\mathrm{on}}$) variability increase $\dot{s}_{\mathrm{mes}}$ for $\overline{\sigma}_{\mathrm{off}}\ll\sigma_{\mathrm{on}}$ ($\sigma_{\mathrm{off}}\gg\overline{\sigma}_{\mathrm{on}}$), but also significantly lower it for most of the space of parameters outside that regime (see Fig. 8(d)). Additionally, in Fig. 8(c) it is shown that $e_{\sigma_{\mathrm{off}}}$ ($e_{\sigma_{\mathrm{on}}}$) is non-monotonic in the variance of the extrinsic noise, and that for small degradation rates there is a finite valued maximum. The summary is that the irreversibility inferred in a single cell can be significantly perturbed by the presence of cell-to-cell variability. This leads to the surprising conclusion that the suppression of cell-to-cell variability is may be more beneficial on some parameters than for others, that is, if the results of the previous section on the observed avoidance of high $\dot{s}_{\mathrm{mes}}$ (i.e., mesoscopic energy expenditure minimization) are to be taken seriously. However, until data is collected on the extrinsic variation of individual kinetic parameters, we cannot be sure whether a population-level EPR provides an evolutionary motivated constraint on the expression noise of individual genes.
Discussion
@@
This study highlights the role of detailed balance and irreversibility in the canonical two-state model of mRNA expression. This includes: the nature of probability fluxes in the steady state (Fig. 3); the analytics of entropy production for a single gene; detailed data analyses of seven kinetic parameter datasets (Figs. 4 and 5 and corresponding figures in the SI); and the counterintuitive effects of cell-to-cell variability on the entropy production rate across a population of cells (Fig. 8). Our results have highlighted a potential mesoscopic analog of the principle of energy expenditure minimization that reveals itself through the EPR of mRNA expression [govern2014optimal]. This finding is aligned with a recent study claiming the non-equilibrium systems reduce their EPR at steady state for large system sizes [ray2025large]. Notably, this is not a thermodynamic phenomenon—we have shown that the lower bound on the EPR predicted by our mesoscopic analysis is not tight compared to order of magnitude estimates. Finally, our results highlight the benefits of transcriptional bursting from the perspective that a burstier behaviors can lead to a lower EPR (Fig. 5(d)).
@@
One of the limitations of our study is that we consider only a two-state gene model of transcription, whereas intricate experiments have shown that transcription in higher-order organisms is best modeled by more than two gene states [suter2011mammalian, bothma2014dynamic, rodriguez2019intrinsic, bartman2019transcriptional, cao2020stochastic, tunnacliffe2020transcriptional, shelansky2024single]. Recent experimental data even reveals that some genes have sub-Poissonian mRNA expression in fission yeast [weidemann2023minimal], a behavior that aligns with earlier theoretical predictions [mcclure1980rate, mitarai2008generation, choubey2015deciphering]. Although the analytic framework introduced in this study can likely be extended to consider multiple gene states, there are no studies that infer transcriptional kinetics across thousands of genes for models more complex than a two-state model [larsson2019genomic, ramskold2024single, sukys2025cell]. Until such datasets are widely accessible, analytic results for the EPR of more complex gene state models cannot be fully utilized. Finally, our framework assumes gamma-distributed extrinsic noise, but heavy-tailed variation (e.g., log-normal) may further alter EPR bounds. Future work could explore this numerically or via approximations, building on studies like [ham2020extrinsic].
Other recent work in the literature echoes our study. Recent research by Gehri et al. [gehri2024entropy] also utilizes a stochastic thermodynamic approach to study the entropy production of models of transcription—in some cases models with three gene states. Therein, they promote the directed information rate as a measure of the throughput in cellular signal processing. Their hypothesis that maximizing the directed information rate is beneficial for transcription is distinct from the data-driven hypothesis of our paper, that most genes actively try not to have a large EPR for mRNA expression. Whereas the focus in our paper is on the EPR of mRNA expression given a two-state promoter, the focus of [gehri2024entropy] is in the EPR resulting from gene state transitions. They hypothesize that energy dissipation limits the information throughput in transcription—i.e., that energy needs to be expended for a greater rate of information transmission. This result is not contradictory to the results presented herein—they could even be complementary, in the case that mRNA expression with an abnormally high EPR could correspond to essential genes that have a greater information throughput.
While we focus on the minimization of the EPR of mRNA expression as a design principle for transcription, other work has focused on the minimization of the coefficient of variation of protein expression as being beneficial—in models of genetic autoregulation [bokes2017gene]. The minimization of noise at the level of mRNA expression would contradict the minimization of mRNA expression EPR due to the observed negative correlation between mRNA expression noise and EPR in real data (see Fig. 4(d) and corresponding figures in the SI). This perhaps hints towards different principles guiding the observed expression of mRNA and proteins.
@@
The finding that cell-to-cell variability can amplify or suppress irreversibility interacts with the seminal work of Lestas et al. [lestas2010fundamental] in several ways. Therein, they investigated the limits of noise suppression given thermodynamic and information theoretic constraints, and derived quantitative bounds on the degree to which noise suppression is actually possible. Their findings, that the suppression of noise in one component of a system may amplify it in another, has clear parallels to the findings presented here, in that accepting a degree of cell-to-cell variability can lead to an ensemble of cells that is more reversible than an equivalent population of identical cells. Where noise is tolerated in the cellular environment is the result of a complex evolutionary process, but this process may not depend only on the properties of a single cell, but on the properties of a population of cells. Our study hints to further evolutionary constraints, that not only is the placement of noise and its suppression important in the context of intrinsic stochasticity, but in where extrinsic noise is tolerated, and in how it is suppressed [weissman2025beyond, weinreich2025population].
@@
In future studies, it would be interesting to study the information theoretic properties of mRNA expression with more complex gene-state architectures—especially if advances in live-seq [chen2022live] and merFISH [zhang2021spatially] data allow for comprehensive datasets of kinetic parameters for more microscopic models of mammalian transcription. Time-resolved mRNA sequencing would also allow for studies into the breaking of detailed balance in dynamic cellular environments, with fluctuating and correlated mRNA expression, using methods introduced in [battle2016broken, lynn2021broken]. Finally, an experimental understanding of how the kinetic parameters of the telegraph model fluctuate across a population of cells could give credence to some of the implications of cell-to-cell variability on reversibility found in this study, and of other studies related to the explicit modeling of extrinsic noise in addition to intrinsic stochasticity [grima2023quantifying, holehouse2020steady, battich2015control, thomas2019intrinsic]. How cells decide where to tolerate cell-to-cell variability is an important open problem.
@@
Acknowledgements: This work was supported by a Lou Schuyler grant from the Santa Fe Institute and National Science Foundation grants DMR-191073 and 2133863. J.H. would like to thank Jacob Calvert, Christopher Lynn, Anish Pandya, Brandon Schlomann, Harrison Hartle, Artemy Kolchinsky and Ramon Grima for discussions and constructive feedback. Special thanks go to Ramon Grima and Artemy Kolchinsky for invaluable suggestions. Finally, J.H. would like to thank an anonymous reviewer for emphasizing the use of statistical null models that serve as a comparison to the real data used in this study.
@@
Data availability: The data for telegraph model parameters used in this paper can be found in [sukys2025cell, ramskold2024single, larsson2019genomic]. The data for the mRNA degradation rates can be found in [sharova2009database].
@@
Author contributions: J.H. conceived, designed, implemented and wrote the study.
@@
In this section, we solve for the eigenvalues of the master equation describing the two-state gene model in Eq. (1). The reaction scheme with $\rho_{\mathrm{off}}=0$ has already been solved in time in the present form [iyer2009stochasticity, wu2023solving] and additionally for the case where mRNA production is bursty [cao2018linear]. Here we specify the calculation for non-zero $\rho_{\mathrm{off}}$. Physically, the eigenvalues of the master equation correspond to the inverse of the fundamental timescales governing the relaxation towards the steady state. We find the eigenvalues by imposing physical constraints on the generating function, essentially that the generating function’s power series is real and non-singular for $z\in[-1,1]$.
@@
Using the separation of variables ansatz $G(x,t)\sim e^{-\lambda_{m}t}f_{m}(x)$, which arises naturally from the linear structure of the master equation, leads to [van1992stochastic, tauber2014critical],
@@
The second solution here is only linearly independent if the exponent is not a integer less than or equal to 0. Now, the condition on $f_{m}(z)$ here necessary to determine the $\lambda_{m}$ is that the powers of (z1)(z-1) pre-multiplying the confluent hypergeometric functions must be integer powers. This is standard practice in eigenfunction solutions which are hypergeometrics, and it means that each $f_{m}(z)$ is real and non-singular for $z\in[-1,1]$, and corresponds to the infinite state limit of other similar solutions found in the literature for Moran-like processes [ewens2004mathematical, holehouse2022exact, mckane2000mean]. Additionally, the exponential pre-factor and the hypergeometric functions themselves are already real and non-singular for finite $z$. Enforcing that $\lambda_{m}=dm$ and $\lambda_{m}-\Sigma=dm$ for integer $m\in\{1,2,3,\ldots\}$ then gives two respective sets of eigenvalues, $\lambda_{m}^{1}=dm$ and $\lambda_{m}^{2}=dm+\Sigma$ that are dependent on the rates of gene switching compared to the degradation rates of the mRNAs.
@@
These eigenvalues represent the inverse of the relaxation time scales of the system, and are in correspondence with results in the fast-switching (marginal single gene state) limit found in [jia2018relaxation]. This almost completes the solution for the generating function of the telegraph model, aside from determination of the constants $C_{m}^{1}$ and $C_{m}^{2}$ which in principle can be done using methods related to Sturm-Liouville theory alongside knowledge of the initial conditions of $G$. For a relevant summary on Sturm-Liouville methods see [holehouse2022exact, Appendix A], for the initial conditions of the telegraph model see the Appendix of [iyer2009stochasticity].
S3 Marginality and coarse-graining of gene expression
A classic form of model reduction is timescale separation, a type of coarse-graining, which is often employed in gene expression under the ‘fast gene switching assumption’, i.e., the condition that $\mathrm{min}\{\sigma_{\mathrm{off}},\sigma_{\mathrm{on}}\}\gg\mathrm{max}\{\rho_{\mathrm{off}},\rho_{\mathrm{on}},d\}$ [holehouse2019revisiting]. Although there are several ways of employing this via quasi equilibrium or steady state assumptions [haseltine2002approximate, rao2003stochastic, kim2015relationship], the most principled way is via the method of ‘averaging’ [bo2017multiple, jia2020dynamical]. Experimental evidence for fast gene switching can be found in [sepulveda2016measurement] among other studies. Denoting $f=\sigma_{\mathrm{off}}/(\sigma_{\mathrm{off}}+\sigma_{\mathrm{on}})$, the method of averaging then gives the following timescale reduced model valid under the fast gene switching assumption,
$G\xrightarrow{\rho}G+M,\;M\xrightarrow{d}\varnothing,$ (S9)
where $\rho=f\rho_{\mathrm{off}}+(1-f)\rho_{\mathrm{on}}$, which is nothing other than a one-dimensional microscopically reversible Markov chain which satisfies detailed balance and hence has zero entropy production in the steady state [gardiner1985handbook]. However, taking the same fast switching limit of Eq. (10) reveals,
which is generally non-zero. This trivial example clearly shows that although common methods of model reduction may capture the dynamics of stochastic gene expression, they lose thermodynamic information, a point made in numerous other studies [jia2016model, bo2014entropy]. Notably, as stated in [bo2014entropy],
A more salient aspect of the analysis presented herein regards the interpretation of the marginal detailed balance without timescale separation arguments. Experiments often only give insight into the state of mRNA or protein numbers, and rarely into the gene states that are theoretically constructed. Hence, if all that can be observed is the mRNA dynamics, the two-state model is equivalent to a birth-death process with a time-dependent mRNA production rate $\rho(t)$, as shown in Fig. S1. Although the equivalent system has a noisy time-dependent rate (whose statistics are specified in the caption), it is still Markovian, and it provides an example of non-zero entropy production in a system that satisfies detailed balance. That is, a case of detailed balance with non-zero EPR. The caveat here is that although the dynamics are Markovian, the time-dependence of $\rho(t)$ allows for this seemingly contradictory situation. When $\rho_{\mathrm{off}}=0$ one can even have detailed balance and an infinite entropy production rate. This provides a striking departure from the standard narrative that detailed balance is reserved for systems in contact with a single reservoir of energy and/or particles or multi-component closed systems [van1976equilibrium, van1992stochastic]. A key takeaway is that although detailed-balance and zero EPR are often stated as being synonymous, this is not necessarily true when the time-independence of the kinetic rates is violated, as can be the case for systems with hidden Markov states.
S4 Calculating the entropy production rate
Following Schnakenberg [schnakenberg1976network, Eqs. (7.6)-(7.11)], the entropy production rate (EPR) is given by,
@@
The first term on the right-hand side is simply $\mathrm{d}S(t)/\mathrm{d}t$, where $S(t)=-\sum_{\mathbf{x}}P(\mathbf{x},t)\ln(P(\mathbf{x},t))$, i.e., the derivative of the entropy change of the system. The second term in the sum is then the entropy change due to coupling with external reservoirs, typically known as the entropy flow—the literature commonly defines this as the negative of the entropy flow rate, since one considers work done by the system, not on the system. In the case where the system is in a non-equilibrium steady state, then the entropy change in the system becomes zero, i.e., $\mathrm{d}S(t)/\mathrm{d}t=0$, and the EPR becomes the negative of the entropy flow rate [esposito2010three, boeger2022energetics, kirchberg2023energy],
@@
Once the connected states have been identified, we arrive at Eq. (9) in the main text. In Eq. (9) there are 2 sums that need to be evaluated. The first is $\sum_{n=1}^{\infty}a_{n}$ which simply evaluates to $b_{1}$ since $a_{n}=b_{n+1}-b_{n}$. One also needs to evaluate the sum $\mathcal{S}=\sum_{n=1}^{\infty}b_{n}$, where, for reference, $b_{n}$ is given by
@@
S5 Extrinsic noise expressions
@@
Figure S2: Analog of Fig. 4 but for 1,436 G2M cell-cycle stage mouse fibroblast genes [sukys2025cell].
Figure S3: Analog of Fig. 4 but for 5,623 mouse fibroblast genes from [ramskold2024single].
Figure S4: Analog of Fig. 4 but for 5,227 C57 mouse embryonic stem cell genes from [larsson2019genomic].
Figure S5: Analog of Fig. 4 but for 4,925 CAST mouse embryonic stem cell genes from [larsson2019genomic].
Figure S6: Analog of Fig. 4 but for 5,858 C57 mouse fibroblast genes from [larsson2019genomic].
Figure S7: Analog of Fig. 4 but for 5,854 CAST mouse fibroblast genes from [larsson2019genomic].
@@
Figure S9: Analog of Fig. 5 but for G2M cell-cycle stage mouse fibroblasts [sukys2025cell].
Figure S10: Analog of Fig. 5 but for mouse fibroblast cells from [ramskold2024single].
Figure S11: Analog of Fig. 5 but for C57 mouse embryonic stem cells from [larsson2019genomic].
Figure S12: Analog of Fig. 5 but for CAST mouse embryonic stem cells from [larsson2019genomic].
Figure S13: Analog of Fig. 5 but for C57 mouse fibroblast cells from [larsson2019genomic].
Figure S14: Analog of Fig. 5 but for CAST mouse fibroblast cells from [larsson2019genomic].
Figure S15: Analog of Fig. 6 but for G2M cell-cycle stage mouse fibroblasts [sukys2025cell].
Figure S16: Analog of Fig. 7 but for G2M cell-cycle stage mouse fibroblasts [sukys2025cell].